OpenStudy (anonymous):

What is the equation of the line that passes through (–2, –3) and is perpendicular to 2x – 3y = 6?

OpenStudy (danjs):

First put the given equation into Y=mX+b form

OpenStudy (danjs):

take the negative reciprical of the slope m to find the Perpendicular Slope

OpenStudy (anonymous):

how do i do that

OpenStudy (danjs):

ok we will walk through it

OpenStudy (anonymous):

ok

OpenStudy (danjs):

2x - 3y = 6 solve for y

OpenStudy (anonymous):

+3y to both sides?

OpenStudy (danjs):

3y = 2x - 6 y = (2/3)x - 2

OpenStudy (danjs):

The slope of that line is 2/3

OpenStudy (anonymous):

how ddid you get from -6 to -2

OpenStudy (danjs):

divide everything by 3

OpenStudy (anonymous):

oh ok

OpenStudy (danjs):

so the slope of the given line is m = 2/3

OpenStudy (anonymous):

but thats none of my answer choics

OpenStudy (danjs):

key point: the slope of a line perpendicular to a given line is flip the slope over and put a negative on it m=-3/2

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

well i found the answer then

OpenStudy (danjs):

now use that slope and the given point (-2,-3) for the line perpendicular

OpenStudy (anonymous):

-3/2x-6

OpenStudy (danjs):

y-y1 = m(x-x1) m=-3/2 (x1,y1) = (-2,-3)

OpenStudy (danjs):

y+3 = (-3/2)*(x+2)

OpenStudy (anonymous):

Thank you mr DanJS

OpenStudy (anonymous):

help me with another one?

OpenStudy (danjs):

y+ 3 = (-3/2)x -3 y = (-3/2)x - 6

OpenStudy (danjs):

sure, just remember a perpendicular slope is found by taking the negative reciprical, if m = 4/5 then the perpendicular slope is -5/4

OpenStudy (danjs):

open a new thread and tag me for another question